# Complex exponential integral means spectrums of univalent functions and the Brennan conjecture

### v6

@@
Abstract
In this paper we investigate the complex exponential integral means spectrums of univalent functions in the unit disk. We show that all integral means spectrum (IMS) functionals for complex exponents on the universal Teichmüller space, the closure of the universal Teichmüller curve, and the universal asymptotic Teichmüller space are continuous. We also show that the complex exponential integral means spectrum of any univalent function admitting a quasiconformal extension to the extended complex plane is strictly less than the universal integral means spectrum. These extend some related results in our recent work [38]. Here we employ a different and more direct approach to prove the continuity of IMS functional on the universal asymptotic Teichmüller space. Additionally, we completely determine the integral means spectrums of all univalent rational functions in the unit disk. As a consequence, we show that the Brennan conjecture is true for this class of univalent functions. Finally, we present some remarks and raise some problems and conjectures regarding IMS functionals on Teichmüller spaces, univalent rational functions, and a multiplier operator whose norm is closely related to the Brennan conjecture.
00footnotetext: 2020 Mathematics Subject Classification. Primary 30C55; Secondary 30C62.00footnotetext: Key words and phrases. Complex exponential integral means spectrum of univalent function, universal integral mean spectrum, Brennan conjecture, univalent function with a quasiconformal extension, univalent rational function, univalent polynomial, universal Teichmüller space, universal asymptotic Teichmüller space, universal Teichmüller curve.00footnotetext: The author was supported by National Natural Science Foundation of China (Grant Nos. 11501157).
1 Introduction
@@
$\beta_{f}(\tau):=\limsup_{r\rightarrow 1^{-}}\frac{\log\int_{-\pi}^{\pi}|[f^{\prime}(re^{i\theta})]^{\tau}|d\theta}{|\log(1-r)|}.$ (1.1)
@@
The universal integral mean spectrums $B(\tau)$ and $B_{b}(\tau)$ are defined as
@@
Because of the important relation to the fine properties of harmonic measures in the complex plane, see [43], the universal integral means spectrums of univalent functions have been studied extensively in recent years. It is an important and difficult problem to find out the exact values of the universal integral mean spectrums $B(\tau)$ and $B_{b}(\tau)$. We shall first review some main known results and open problems for real $\tau$ in this topic. It was first observed by Makarov in [43] that
@@
There have been some studies on the integral means spectrums of univalent functions admitting a quasiconformal extension, see for example [33], [34], [35], [36], [37], [49] and [50]. From Theorem 1.9, we can see the problem that determining the exact values of $B_{b}(\tau)$ as a global extremal problem in the class $\mathcal{S}_{q}$. To enrich the understanding of the theory of integral means spectrums of univalent functions, in the present paper, we consider the following problem, which differs from previous studies that finding a better estimation for $B_{b}(\tau)$.
@@
The rest of paper is organized as follows. In the next section, we will fix some notations used in this paper and recall some basic properties about univalent functions, quasiconformal mappings, and the definitions of quasiconformal Teichmüller spaces. Also, Theorem 1-4 will be restated in Section 2. We will provide some lemmas in Section 3. We shall give the proof of Theorem 1 and 2 in Section 4. We will prove Theorem 3 and 4 in Section 5. In Section 6, we study the integral means spectrums of univalent rational functions in $\Delta$ and prove that the Brennan conjecture is true for this class of functions. In Section 7, for IMS functionals on Teichmüller spaces, univalent rational functions, and a multiplier operator whose norm is closely related to the Brennan conjecture, we finally present some remarks and raise some problems and conjectures.
2 Preliminaries and restatement of the first main results
@@
2.1 Univalent functions and quasiconformal mappings
@@
$N_{g\circ f}(z)=N_{g}(f(z))[f^{\prime}(z)]+N_{f}(z),\,\,z\in\Omega,$ (2.1)
@@
$S_{g\circ f}(z)=S_{g}(f(z))[f^{\prime}(z)]^{2}+S_{f}(z),\,\,z\in\Omega.$ (2.2)
@@
$\mu_{g\circ f^{-1}}\circ f(z)=\frac{1}{\chi}\frac{\mu_{g}(z)-\mu_{f}(z)}{1-\overline{\mu_{f}(z)}\mu_{g}(z)},\,\,\chi=\frac{\overline{\partial f}}{\partial f},\,\,\,z\in\Omega_{1}.$ (2.3)
@@
$b(f):=\inf\{\|\mu_{f}|_{\Omega^{*}-E}\|_{\infty}:\,E{\text{ is a compact set in}}\,\,\Omega^{*}\}.$ (2.4)
@@
2.2 Quasiconformal Teichmüller spaces
@@
$h^{*}(\mu)=\inf\{\|\mu|_{\Omega^{*}-E}\|_{\infty}:\,E{\text{ is a compact set in}}\,\,\Omega^{*}\}.$ (2.5)
@@
Note that for any $f\in\mathcal{S}_{q}$, we can find a $\mu\in M(\Delta^{*})$ such that $f=\sigma\circ f_{\mu}$, $\sigma$ is a Möbius transformation. Then we check from Theorem 1.9 that $B_{b}(\tau)=\sup\limits_{{[\mu]_{T}}\in T}\beta_{f_{\mu}}(\tau)$ for each $\tau\in\mathbb{C}$.
2.3 Restatement of the first main results
@@
$f_{\phi}(z)=\int_{0}^{z}e^{\int_{0}^{\zeta}\phi(w)dw}d\zeta,\,\,\,z\in\Delta.$ (2.6)
@@ 3 Five lemmas
$|N_{\mathbf{h}}(\zeta)|\rho_{\Omega}^{-1}(\zeta)\leq 8\|\mu_{\mathbf{h}}\|_{\infty},$ (3.1)
@@
$\log[{\mathbf{h}}^{\prime}\circ f(z)]^{\tau}=\tau\log H(z)=\tau\int_{z_{0}}^{z}G(\zeta)d\zeta+\tau\log H(z_{0}),$ (3.2)
@@
$C_{1}(r_{0},\tau,\varepsilon)\Big(\frac{1-|z|}{1+|z|}\Big)^{\frac{\varepsilon}{2}}\leq|{\mathbf{h}}^{\prime}\circ f(z)|\leq C_{2}(r_{0},\tau,\varepsilon)\Big(\frac{1+|z|}{1-|z|}\Big)^{\frac{\varepsilon}{2}},$ (3.4)
@@
$\iint_{\Delta}|[f^{\prime}(z)]^{\tau}|(1-|z|^{2})^{-1+\beta+\varepsilon/2}dxdy<\infty,$ (3.5)
@@
$\iint_{\Delta}|[f^{\prime}(z)]^{\tau}|(1-|z|^{2})^{-1+\beta-\varepsilon/2}dxdy=\infty.$ (3.6)
@@ 4 Proof of Theorem 1 and 2
4.1 Isomorphism mapping between Teichmüller spaces
@@
$\left|\sigma(\mu,\nu)(z)\right|=\left|\frac{\mu^{\sharp}-\nu^{\sharp}}{1-{\overline{\nu^{\sharp}}\mu^{\sharp}}}\circ f_{\omega}(z)\right|=\left|\sigma(\mu^{\sharp},\nu^{\sharp})(\zeta)\right|,\,\zeta=f_{\omega}(z),\,z\in\Delta^{*}.$ (4.1)
@@
$\left|\sigma(\mu,\nu)(\zeta)\right|=\left|\frac{{}^{\sharp}\mu-{{}^{\sharp}}\nu}{1-{\overline{{}^{\sharp}\nu}{{}^{\sharp}}\mu}}\circ f_{\omega}^{-1}(\zeta)\right|=\sigma({{}^{\sharp}}\mu,{{}^{\sharp}}\nu)(z),\,z\in\Delta^{*},\,\zeta=f_{\omega}(z).$ (4.2)
@@
4.2 Extremal theory of quasiconformal mappings
@@ 4.3 A claim
$d_{AT}([\mu]_{AT(\Omega)},[\nu]_{AT(\Omega)})$ $=$ $\inf\limits_{\mu_{1}\approx\mu,\nu_{1}\approx\nu}\Big\{\tanh^{-1}\|h^{*}(\sigma(\mu_{1},\nu_{1}))\|_{\infty}\Big\}$ (4.3)
@@ 4.4 Proof of Theorem 2
In particular, $\beta_{f_{\mu}}(\tau)=0$ for any $\tau\in\mathbb{C}$ if $\mu\approx 0$ in $M(\Delta^{*})$. We call $f_{\mu}$ is an asymptotically conformal mapping in $\Delta$ when $\mu\approx 0$ in $M(\Delta^{*})$. In general, we say a function $f\in\mathcal{S}_{b}$ is an asymptotically conformal mapping in $\Delta$ if it admits a quasiconformal extension (still denoted by $f$) to $\widehat{\mathbb{C}}$ with $\mu_{f}(z)\rightarrow 0$ as $|z|\rightarrow 1^{+}$, see [25], [46], [47]. From Proposition 4.7, we see that $B_{b}(\tau)=\sup\limits_{[\mu]_{AT}\in AT}\beta_{f_{\mu}}(\tau)$ for each $\tau\in\mathbb{C}$.
@@
5 Proof of Theorem 3 and 4
5.1 Proof of Theorem 3
@@ 5.2 Proof of Theorem 4
$\int_{-\pi}^{\pi}|[f^{\prime}({\bf r}_{n}e^{i\theta})]^{\tau_{0}+t\tau_{0}}|d\theta$ $>$ $2^{-1-t}\pi^{-t}{(1-{\bf r}_{n})^{-(\beta-\varepsilon)(1+t)}}.$ (5.1)
@@
$\int_{-\pi}^{\pi}|[f^{\prime}({\bf r}_{n}e^{i\theta})]^{\tau_{0}+t\tau_{0}}|d\theta$ $>$ $16^{-1-t}\pi^{-t}{(1-{\bf r}_{n})^{-(\beta-\varepsilon)(1+t)}}.$ (5.2)
@@
$\int_{-\pi}^{\pi}|[f^{\prime}({\bf r}_{n}e^{i\theta})]^{\tau_{0}+t\tau_{0}}|d\theta$ $>$ $16^{-1-t}\pi^{-t}{(1-{\bf r}_{n})^{-(\beta-\varepsilon)(1+t)}}.$ (5.3)
@@
$\beta_{f}(\tau_{0}+t\tau_{0})\geq\beta+\frac{1}{3}\beta t>\beta=\beta_{f}(\tau_{0}).$
@@
$\mathbf{h}^{\prime}\circ f(z)=[f^{\prime}(z)]^{\varepsilon},z\in\Delta,$ (5.4)
@@
$\mathbf{h}^{\prime\prime}\circ f(z)\cdot f^{\prime}(z)=\varepsilon[f^{\prime}(z)]^{\varepsilon-1}f^{\prime\prime}(z),z\in\Delta.$ (5.5)
@@
6 Integral means spectrums of the univalent rational functions
In this section, we study the integral means spectrums of univalent rational functions in $\Delta$. In the rest of the paper, the notation $P_{n}(z)$ denotes a polynomial of degree at most $n$, $n\in\mathbb{N}\cup\{0\}$. In particular, $P_{n}$ is a complex constant when $n=0$. We denote by $\mathcal{R}$ the class of all rational functions $R$ in $\Delta$ with $R(0)=R^{\prime}(0)-1=0$ and which have no poles in $\Delta$. Let $\Omega$ be a subset of the complex plane, we say $R\in\mathcal{R}$ has no zeros, or one zero in $\Omega$ if the equation $R(z)=0$ has no roots, or one root in $\Omega$, respectively. We will say $R\in\mathcal{R}$ has no critical points, or one critical point if the equation $R^{\prime}(z)=0$ has no roots, or one root in $\Omega$, respectively. A rational function $R\in\mathcal{R}$ can be uniquely represented as
@@ 6 Integral means spectrums of the univalent rational functions
6.1 Univalent rational functions and the main result of this section

### v7

# Complex exponential integral means spectra of univalent functions and Brennan conjecture

@@
Abstract.
In this paper we investigate the complex exponential integral means spectra of univalent functions in the unit disk. We show that all integral means spectrum (IMS) functionals for complex exponents on the universal Teichmüller space, the closure of the universal Teichmüller curve, and the universal asymptotic Teichmüller space are continuous. We also show that the complex exponential integral means spectrum of any univalent function admitting a quasiconformal extension to the extended complex plane is strictly less than the universal integral means spectrum. These results extend some related ones in recent work [38] by Jin. Here we employ a different and more direct approach to prove the continuity of IMS functional on the universal asymptotic Teichmüller space. Additionally, we completely determine the integral means spectra of all univalent rational functions in the unit disk. As a consequence, we show that the Brennan conjecture is true for this class of univalent functions. Finally, we present some remarks and raise some problems and conjectures regarding IMS functionals on Teichmüller spaces, univalent rational functions, and a multiplication operator whose norm is closely related to the Brennan conjecture.
Key words and phrases:
Complex exponential integral means spectrum of univalent function, universal integral mean spectrum, Brennan conjecture, univalent function with a quasiconformal extension, univalent rational function, univalent polynomial, universal Teichmüller space, universal asymptotic Teichmüller space, universal Teichmüller curve
2020 Mathematics Subject Classification:
Primary 30C55; Secondary 30C62
The author was supported by National Natural Science Foundation of China (Grant Nos. 11501157).
1. Introduction
@@
(1.1) $\beta_{f}(\tau):=\limsup_{r\rightarrow 1^{-}}\frac{\log\int_{-\pi}^{\pi}|[f^{\prime}(re^{i\theta})]^{\tau}|d\theta}{|\log(1-r)|}.$
@@
The universal integral mean spectra $B(\tau)$ and $B_{b}(\tau)$ are defined as
@@
Because of the important relation to the fine properties of harmonic measures in the complex plane, see [43], the universal integral means spectra of univalent functions have been studied extensively in recent years. It is an important and difficult problem to find out the exact values of the universal integral mean spectra $B(\tau)$ and $B_{b}(\tau)$. We shall first review some main known results and open problems for real $\tau$ in this topic. It was first observed by Makarov in [43] that
@@
There have been some studies on the integral means spectra of univalent functions admitting a quasiconformal extension, see for example [33], [34], [35], [36], [37], [49] and [50]. From Theorem 1.9, we can see the problem that determining the exact values of $B_{b}(\tau)$ as a global extremal problem in the class $\mathcal{S}_{q}$. To enrich the understanding of the theory of integral means spectra of univalent functions, in the present paper, we consider the following problem, which differs from previous studies that finding a better estimation for $B_{b}(\tau)$.
@@
The rest of paper is organized as follows. In the next section, we will fix some notations used in this paper and recall some basic properties about univalent functions, quasiconformal mappings, and the definitions of quasiconformal Teichmüller spaces. Also, Theorem 1-4 will be restated in Section 2. We will provide some lemmas in Section 3. We shall give the proof of Theorem 1 and 2 in Section 4. We will prove Theorem 3 and 4 in Section 5. In Section 6, we study the integral means spectra of univalent rational functions in $\Delta$ and prove that the Brennan conjecture is true for this class of functions. In Section 7, for IMS functionals on Teichmüller spaces, univalent rational functions, and a multiplication operator whose norm is closely related to the Brennan conjecture, we finally present some remarks and raise some problems and conjectures.
2. Preliminaries and restatement of the first main results
@@
2.1. Univalent functions and quasiconformal mappings
@@
(2.1) $N_{g\circ f}(z)=N_{g}(f(z))[f^{\prime}(z)]+N_{f}(z),\,\,z\in\Omega,$
@@
(2.2) $S_{g\circ f}(z)=S_{g}(f(z))[f^{\prime}(z)]^{2}+S_{f}(z),\,\,z\in\Omega.$
@@
(2.3) $\mu_{g\circ f^{-1}}\circ f(z)=\frac{1}{\chi}\frac{\mu_{g}(z)-\mu_{f}(z)}{1-\overline{\mu_{f}(z)}\mu_{g}(z)},\,\,\chi=\frac{\overline{\partial f}}{\partial f},\,\,\,z\in\Omega_{1}.$
@@
(2.4) $b(f):=\inf\{\|\mu_{f}|_{\Omega^{*}-E}\|_{\infty}:\,E{\text{ is a compact set in}}\,\,\Omega^{*}\}.$
@@
2.2. Quasiconformal Teichmüller spaces
@@
(2.5) $h^{*}(\mu)=\inf\{\|\mu|_{\Omega^{*}-E}\|_{\infty}:\,E{\text{ is a compact set in}}\,\,\Omega^{*}\}.$
@@
Note that for any $f\in\mathcal{S}_{q}$, we can find a $\mu\in M(\Delta^{*})$ such that $f=\sigma\circ f_{\mu}$, $\sigma$ is a Möbius transformation. Then we check from Theorem 1.9 that
$B_{b}(\tau)=\sup\limits_{{[\mu]_{T}}\in T}\beta_{f_{\mu}}(\tau)$
for each $\tau\in\mathbb{C}$.
2.3. Restatement of the first main results
@@
(2.6) $f_{\phi}(z)=\int_{0}^{z}e^{\int_{0}^{\zeta}\phi(w)dw}d\zeta,\,\,\,z\in\Delta.$
@@ 3. Five lemmas
(3.1) $|N_{\mathbf{h}}(\zeta)|\rho_{\Omega}^{-1}(\zeta)\leq 8\|\mu_{\mathbf{h}}\|_{\infty},$
@@
(3.2) $\log[{\mathbf{h}}^{\prime}\circ f(z)]^{\tau}=\tau\log H(z)=\tau\int_{z_{0}}^{z}G(\zeta)d\zeta+\tau\log H(z_{0}),$
@@
(3.4) $C_{1}(r_{0},\tau,\varepsilon)\Big(\frac{1-|z|}{1+|z|}\Big)^{\frac{\varepsilon}{2}}\leq|{\mathbf{h}}^{\prime}\circ f(z)|\leq C_{2}(r_{0},\tau,\varepsilon)\Big(\frac{1+|z|}{1-|z|}\Big)^{\frac{\varepsilon}{2}},$
@@
(3.5) $\iint_{\Delta}|[f^{\prime}(z)]^{\tau}|(1-|z|^{2})^{-1+\beta+\varepsilon/2}dxdy<\infty,$
@@
(3.6) $\iint_{\Delta}|[f^{\prime}(z)]^{\tau}|(1-|z|^{2})^{-1+\beta-\varepsilon/2}dxdy=\infty.$
@@ 4. Proof of Theorem 1 and 2
4.1. Isomorphism mapping between Teichmüller spaces
@@
(4.1) $\left|\sigma(\mu,\nu)(z)\right|=\left|\frac{\mu^{\sharp}-\nu^{\sharp}}{1-{\overline{\nu^{\sharp}}\mu^{\sharp}}}\circ f_{\omega}(z)\right|=\left|\sigma(\mu^{\sharp},\nu^{\sharp})(\zeta)\right|,\,\zeta=f_{\omega}(z),\,z\in\Delta^{*}.$
@@
(4.2) $\left|\sigma(\mu,\nu)(\zeta)\right|=\left|\frac{{}^{\sharp}\mu-{{}^{\sharp}}\nu}{1-{\overline{{}^{\sharp}\nu}{{}^{\sharp}}\mu}}\circ f_{\omega}^{-1}(\zeta)\right|=\sigma({{}^{\sharp}}\mu,{{}^{\sharp}}\nu)(z),\,z\in\Delta^{*},\,\zeta=f_{\omega}(z).$
@@
4.2. Extremal theory of quasiconformal mappings
@@ 4.3. A claim
(4.3) $d_{AT}([\mu]_{AT(\Omega)},[\nu]_{AT(\Omega)})$ $=$ $\inf\limits_{\mu_{1}\approx\mu,\nu_{1}\approx\nu}\Big\{\tanh^{-1}\|h^{*}(\sigma(\mu_{1},\nu_{1}))\|_{\infty}\Big\}$
@@ 4.4. Proof of Theorem 2
In particular, $\beta_{f_{\mu}}(\tau)=0$ for any $\tau\in\mathbb{C}$ if $\mu\approx 0$ in $M(\Delta^{*})$. We call $f_{\mu}$ is an asymptotically conformal mapping in $\Delta$ when $\mu\approx 0$ in $M(\Delta^{*})$. In general, we say a function $f\in\mathcal{S}_{b}$ is an asymptotically conformal mapping in $\Delta$ if it admits a quasiconformal extension (still denoted by $f$) to $\widehat{\mathbb{C}}$ with $\mu_{f}(z)\rightarrow 0$ as $|z|\rightarrow 1^{+}$, see [25], [46], [47]. From Proposition 4.7, we see that
$B_{b}(\tau)=\sup\limits_{[\mu]_{AT}\in AT}\beta_{f_{\mu}}(\tau)$
for each $\tau\in\mathbb{C}$.
@@
5. Proof of Theorem 3 and 4
5.1. Proof of Theorem 3
@@ 5.2. Proof of Theorem 4
(5.1) $\int_{-\pi}^{\pi}|[f^{\prime}({\bf r}_{n}e^{i\theta})]^{\tau_{0}+t\tau_{0}}|d\theta$ $>$ $2^{-1-t}\pi^{-t}{(1-{\bf r}_{n})^{-(\beta-\varepsilon)(1+t)}}.$
@@
(5.2) $\int_{-\pi}^{\pi}|[f^{\prime}({\bf r}_{n}e^{i\theta})]^{\tau_{0}+t\tau_{0}}|d\theta$ $>$ $16^{-1-t}\pi^{-t}{(1-{\bf r}_{n})^{-(\beta-\varepsilon)(1+t)}}.$
@@
(5.3) $\int_{-\pi}^{\pi}|[f^{\prime}({\bf r}_{n}e^{i\theta})]^{\tau_{0}+t\tau_{0}}|d\theta$ $>$ $16^{-1-t}\pi^{-t}{(1-{\bf r}_{n})^{-(\beta-\varepsilon)(1+t)}}.$
@@
$\beta_{f}(\tau_{0}+t\tau_{0})\geq\beta+\frac{1}{2}\beta t>\beta=\beta_{f}(\tau_{0}).$
@@
(5.4) $\mathbf{h}^{\prime}\circ f(z)=[f^{\prime}(z)]^{\varepsilon},z\in\Delta,$
@@
(5.5) $\mathbf{h}^{\prime\prime}\circ f(z)\cdot f^{\prime}(z)=\varepsilon[f^{\prime}(z)]^{\varepsilon-1}f^{\prime\prime}(z),z\in\Delta.$
@@
6. Integral means spectra of the univalent rational functions
In this section, we study the integral means spectra of univalent rational functions in $\Delta$. In the rest of the paper, the notation $P_{n}(z)$ denotes a polynomial of degree at most $n$, $n\in\mathbb{N}\cup\{0\}$. In particular, $P_{n}$ is a complex constant when $n=0$. We denote by $\mathcal{R}$ the class of all rational functions $R$ in $\Delta$ with $R(0)=R^{\prime}(0)-1=0$ and which have no poles in $\Delta$. Let $\Omega$ be a subset of the complex plane, we say $R\in\mathcal{R}$ has no zeros, or one zero in $\Omega$ if the equation $R(z)=0$ has no roots, or one root in $\Omega$, respectively. We will say $R\in\mathcal{R}$ has no critical points, or one critical point if the equation $R^{\prime}(z)=0$ has no roots, or one root in $\Omega$, respectively. A rational function $R\in\mathcal{R}$ can be uniquely represented as
@@ 6. Integral means spectra of the univalent rational functions
6.1. Univalent rational functions and the main result of this section